# Modeling lower-truncated and right-censored insurance claims with an extension of the MBBEFD class

### v1

@@ Abstract
In general insurance, claims are often lower-truncated and right-censored because insurance contracts may involve deductibles and maximal covers. Most classical statistical models are not (directly) suited to model lower-truncated and right-censored claims. A surprisingly flexible family of distributions that can cope with lower-truncated and right-censored claims is the class of MBBEFD distributions that originally has been introduced by Bernegger (1997) for reinsurance pricing, but which has not gained much attention outside the reinsurance literature. We derive properties of the class of MBBEFD distributions, and we extend it to a bigger family of distribution functions suitable for modeling lower-truncated and right-censored claims. Interestingly, in general insurance, we mainly rely on unimodal skewed densities, whereas the reinsurance literature typically proposes monotonically decreasing densities within the MBBEFD class.
Keywords. General insurance claims, lower-truncation, right-censoring, MBBEFD distribution, unimodal density, skewed density, normalized loss, exposure curve, Swiss Re exposure curve, Lloyd’s exposure curve.
@@ 1 Introduction
Insurance contracts in general insurance often involve deductibles $d>0$ and maximal covers $M>0$. Deductibles are introduced to reduce the number of small claims which mainly cause administrative expenses but which are not essential in risk mitigation. Maximal covers are introduced to control the maximal loss of an insurer. A maximal cover may, e.g., refer to the property value insured (after subtracting the deductible), or to the maximal insurance coverage warranted to a liability claim. Denote by $Y^{\dagger}$ the total financial loss. The insurance claim $Y$ after subtracting the deductible $d>0$ and with a maximal cover of size $M>0$ is given by
$Y=\min\left\{(Y^{\dagger}-d)_{+},\,M\right\}.$ (1.1)
We say, this financial loss is lower-truncated at $d>0$ and right-censored at $M>0$ (after subtracting the deductible). Statistical modeling of lower-truncated and right-censored claims is a notoriously difficult problem because most statistical models have an unbounded support, e.g., the supports of the gamma and the log-normal distributions are the entire positive real line ${\mathbb{R}}_{+}$. In most cases, this implies that fitting a statistical model to lower-truncated and right-censored data is not a problem that is easily analytically tractable. We give an example. We start from a classical statistical model such as the gamma distribution for the total financial loss $Y^{\dagger}\sim F_{\Gamma}$, where $F_{\Gamma}$ denotes the gamma distribution with corresponding gamma density $f_{\Gamma}$ on ${\mathbb{R}}_{+}$. Lower-truncation and right-censoring introduces two difficulties which are illustrated in Figure 1. First, the lower-truncation $(Y^{\dagger}-d)_{+}$ of the total financial loss $Y^{\dagger}$ implies, in general, that the density of the lower-truncated claim is positive in 0, see Figure 1. In the above mentioned gamma case, this means that the lower-truncation with $d>0$ leads to a new density given by
@@
Second, right-censoring at $M>0$ of this lower-truncated claim $(Y^{\dagger}-d)_{+}$ leads to a point mass in $M$, resulting in the density $f$ of the lower-truncated and right-censored claim $Y$
@@
assuming that the response variable $Y$ is absolutely continuous on $(0,M)$ with density $f_{\theta}(y)$, having a point mass $1-F_{\theta}(M_{-})=1-\lim_{y\uparrow M}F_{\theta}(y)$ in $M$, and with (unknown) model parameter $\theta$. Fitting such a model with MLE can be difficult because we need an analytically tractable form for both the density $f_{\theta}(\cdot)$ and its distribution function $F_{\theta}(\cdot)$, see (1.3). This is not the case, e.g., in the lower-truncated and right-censored gamma model given in (1.2). Therefore, in such cases, one either needs to rely on numerical integration of the density (which can be computationally demanding) or one uses a version of the Expectation-Maximization (EM) algorithm by interpreting the lower-truncation and right-censoring as a missing information problem; we refer to Verbelen et al. [16], Fung et al. [5] and Sections 6.4.2 and 6.4.3 in Wüthrich–Merz [17]. However, also this EM algorithm approach has its drawbacks as it requires tractability of conditional tail expectations and reasonable dispersion estimates in multi-dimensional parameter settings. These two side constraints lead to further restrictions on the class of solvable models, e.g., these problems can only be solved for a very small number of models within the class of Tweedie’s models [15], namely, for the Tweedie’s models stated in Theorem 3 of Blæsild–Jensen [3]; we also refer to Landsman–Valdez [6]. In a series of papers, Poudyal [9, 10] and Poudyal–Brazauskas [11] consider trimmed and/or winsorized method of moments estimators for truncated and/or censored data; in statistics, truncation is also called trimming and censoring winsorizing. In these papers, trimming and winsorizing is also shown to be a useful method of robustifying moment estimation under extreme claims.
@@
The goal of this paper is first to study the properties of these MBBEFD densities and to extend it to a bigger class of models that will be called the Bernegger class. Our contribution is to show that the Bernegger class is rather rich including monotonically decreasing densities, unimodal densities and monotonically increasing densities, and our extension provides new families of lower-truncated and right-censored random variables that allow for skewness in the absolutely continuous part of the distribution. This is of particular interest because unimodal skewed densities are suited for modeling lower-truncated and right-censored claims in general insurance since their empirical density roughly looks like the one given in Figure 1. In particular, one can recognize the distribution of right-censored exponential random variables as a special case of the Bernegger class.
@@
Surprisingly, the class of MBBEFD densities of Bernegger [2] has only entered the reinsurance literature; see, e.g., Parodi–Watson [8], Abramson [1], Riegel [13], Chapter 21 of Parodi [7], and the R [12] package mbbefd of Dutang et al. [4, 14]. Popular examples in reinsurance pricing are the so-called Swiss Re and Lloyd’s exposure curves that are special cases of MBBEFD exposure curves; see Bernegger [2]. However, in this reinsurance pricing literature, one mainly focuses on exposure curves and not on the resulting densities nor on their properties. We show that the most popular choices from reinsurance lead to monotonically decreasing densities, whereas we are mainly interested into the unimodal case, as this is the common situation in general insurance pricing, e.g., in private insurance lines.
@@
Organization. This manuscript is organized as follows. In the next section, we state the necessary properties that any exposure curve has to fulfill in order to describe distribution functions allowing to model lower-truncated and right-censored claims. In Section 3, we start from the MBBEFD class of distributions of Bernegger [2] by stating some of its properties, and then we extend it to a richer family of distributions, the Bernegger class. In Section 4, we introduce a subclass of the Bernegger class that incorporates the MBBEFD class of distributions, whereas in Section 5, we treat another subclass of distributions that includes the exponential distribution. In Section 6, we use a real dataset of lower-truncated and right-censored claims in order to compare the performance of the MBBEFD class of distributions to four other classes of distributions when trying to fit those models using maximum likelihood estimation (MLE). Finally, the last section concludes this work. All mathematical proofs are provided in the appendix.
@@ 2.1 From Exposure curves to distributions
In reinsurance claims modeling, one often works with exposure curves instead of distribution functions. Assume we have a positively supported response variable $Y\sim F_{Y}$, and assume that the maximal possible loss (MPL) is given by $M>0$, i.e., $0<Y\leq M$, ${\mathbb{P}}$-a.s. We define the normalized loss $Z=Y/M$. Denote the distribution function of the normalized loss $Z$ by $F_{Z}$, being supported in $(0,1]$. The exposure curve of a normalized loss $Z\sim F_{Z}$ is given by
@@
for $z\in[0,1]$; see Bernegger [2]. This exposure curve $G:[0,1]\to[0,1]$ is non-decreasing, concave, and satisfies the property $G^{\prime}(0)>0$, as well as the normalization $G(0)=0$ and $G(1)=1$; for examples see Figure 2 (lhs), below. In the last integral, a change of variable $s\in[0,1]\mapsto t=sM\in[0,M]$ gives us
@@
the latter being the exposure curve of $Y$ on $[0,M]$. Thus, we can equally work with the responses $Y$ and $Z$, but the normalized losses $Z$ will have the advantage that they live on a common unit interval $[0,1]$. Now, let us take the opposite view and characterize the distribution function of a random variable $Z\sim F_{Z}$ obtained from a function $G:[0,1]\rightarrow\mathbb{R}$ satisfying the same properties of an exposure curve. Under the assumptions of the next theorem, this distribution $F_{Z}$ leads to an absolutely continuous density on $[0,1)$ and a point mass at 1.
@@
for $z\in[0,1)$, and a point mass at 1 given by
@@
As seen previously, if we start from any such function $G$, we can derive a distribution whose density is absolutely continuous and of closed form on $[0,1)$, with a point mass in 1 and a mean that are of closed form too, i.e., we have a class of models that has a full tractable mean, density and point mass, which is suitable to model right-censored claims, and if $G^{\prime\prime}(0)<0$, we can also consider lower-truncation, see (2.2). It turns out that a linear combination of exposure curves allows us to define a mixture of their respective associated distribution functions as the next result shows.
@@
Let $(\alpha_{i})_{i=1}^{n}$ be non-negative weights adding up to 1 and let $(G_{i})_{i=1}^{n}$ be exposure curves leading to distributions functions $(F_{i})_{i=1}^{n}$, densities $(f_{i})_{i=1}^{n}$, and point masses in 1 equal to $(p_{i})_{i=1}^{n}$ as in Theorem 2.1. The convex combination
@@
for $z\in[0,1]$, is again an exposure curve, allowing to define the distribution function of a random variable $Z\sim F_{Z}$ given by
@@ 2.2 Flexibility of the point mass in the right-censoring point
The point mass $p$ in 1 is automatically determined by formula (2.3). Often, one may require more modeling flexibility in the choice of this point mass, while still retaining the tractability of the density and the mean as it was shown in Theorem 2.1. A simple way to do so connects to so-called one-inflated distributions; we refer to Dutang et al. [4]. In the case of Theorem 2.1, this can easily be achieved. The next corollary shows, how we can obtain them by looking at the conditional density of the random variable $Z_{0}\stackrel{{\scriptstyle\rm(d)}}{{=}}Z|_{\{Z<1\}}$.
@@
Let $G:[0,1]\rightarrow\mathbb{R}$ be an exposure curve, we receive an absolutely continuous density on $[0,1)$ given by
@@
which integrates to 1. This density provides the mean for the random variable $Z_{0}\sim f_{0}$
@@
We can now add a point mass $q\in(0,1)$ at 1 to this density. This adds one more parameter to the model, giving us a mixture distribution between an absolutely continuous part $f_{0}$ on $[0,1)$ and a point mass in 1. We have the following corollary.
@@
Note that the transformation achieved in Corollary 2.5 can also be obtained using Lemma 2.3. Indeed, let us consider the function
@@
for $z\in[0,1]$, $w\in[0,1]$, and where $G$ is an exposure curve. Then, the function $\tilde{G}$ is an exposure curve. Using Theorem 2.1, if we denote by $F$ the distribution function obtained from the exposure curve $G$, by $f$ the absolutely continuous density on $[0,1)$ and by $p$ the point mass in 1, we can characterize the distribution function $\tilde{F}$, the absolutely continuous density on $[0,1)$ $\tilde{f}$, and the point mass $\tilde{p}$ obtained from the exposure curve $\tilde{G}$ using
@@
In what follows, the goal will be to introduce examples of exposure curves that are useful to model lower-truncated and right-censored losses. Before that, we start by studying the explicit family of exposure curves introduced by Bernegger [2].
@@ 3.1 The class of MBBEFD exposure curves and densities
The MBBEFD class of Bernegger [2] selects an explicit family of exposure curves. This family is parametrized through two parameters $g\geq 1$ and $b\geq 0$, and is given as follows for $z\in[0,1]$
@@
${\mathbb{E}}_{b,g}[Z]=\left\{\begin{array}[]{ll}\frac{\log(g)}{g-1}&\text{ for$g>1$ and $b=1$,}\\ \frac{b-1}{\log(b)}&\text{ for $g>1$ and $bg=1$,}\\ \frac{b-1}{\log(b)}\frac{\log(bg)}{bg-1}&\text{ for $g>1$, $b>0$, $b\neq 1$ and $bg\neq 1$.}\end{array}\right.$ (3.5)
@@
For $c=1.5,2,3,4$, one obtains the Swiss Re exposure curves, and for $c=5$, the Lloyd’s exposure curve, these are illustrated in Figure 2 (lhs). The right-hand side of this figure shows the resulting densities $f_{b(c),g(c)}(\cdot)$, and we remark that all of the resulting MBBEFD densities are monotonically decreasing on $(0,1)$. $\blacksquare$

### v2

@@ Abstract
In general insurance, claims are often lower-truncated and right-censored because insurance contracts may involve deductibles and maximal covers. Most classical statistical models are not (directly) suited to model lower-truncated and right-censored claims. A surprisingly flexible family of distributions that can cope with lower-truncated and right-censored claims is the class of MBBEFD distributions that originally has been introduced by Bernegger (1997) for reinsurance pricing, but which has not gained much attention outside the reinsurance literature. Interestingly, in general insurance, we mainly rely on unimodal skewed densities, whereas the reinsurance literature typically proposes monotonically decreasing densities within the MBBEFD class. We show that this class contains both types of densities, and we extend it to a bigger family of distribution functions suitable for modeling lower-truncated and right-censored claims. In addition, we discuss how changes in the deductible or the maximal cover affect the chosen distributions.
Keywords. General insurance claims, deductible, maximal cover, lower-truncation, right-censoring, MBBEFD distribution, unimodal density, skewed density, normalized loss, exposure curve, Swiss Re exposure curve, Lloyd’s exposure curve.
@@ 1 Introduction
Insurance contracts in general insurance often involve deductibles $d>0$ and maximal covers $M>0$. Deductibles are introduced to reduce the number of small claims which mainly cause administrative expenses but which are not essential in risk mitigation. Maximal covers are introduced to control the maximal loss of an insurer. A maximal cover may, e.g., refer to the property value insured (after subtracting the deductible), or to the maximal insurance coverage warranted to a liability claim. Denote by $X$ the total financial loss. The insurance claim $Y$ after subtracting the deductible $d>0$ and with a maximal cover of size $M>0$ is given by
$Y=\min\left\{(X-d)_{+},\,M\right\}\,|\,X>d.$ (1.1)
We say, this financial loss is lower-truncated at $d>0$ and right-censored at $M>0$ (after subtracting the deductible). Statistical modeling of lower-truncated and right-censored claims is a notoriously difficult problem. Most statistical models have an unbounded support, e.g., the supports of the gamma and the log-normal distributions are the entire positive real line ${\mathbb{R}}_{+}$. In many cases, this implies that fitting a statistical model to lower-truncated and right-censored data is not a problem that is easily analytically tractable. We give an example. We start from a classical statistical model such as the gamma distribution for the total financial loss $X\sim F_{\Gamma}$, where $F_{\Gamma}$ denotes the gamma distribution with corresponding gamma density $f_{\Gamma}$ on ${\mathbb{R}}_{+}$. Lower-truncation and right-censoring introduces two difficulties which are illustrated in Figure 1. First, the lower-truncation $(X-d)_{+}\,|\,X>d$ of the total financial loss $X$ implies, in general, that the density of the lower-truncated claim is positive in 0, see Figure 1. In the above mentioned gamma case, this means that the lower-truncation with $d>0$ leads to a new density given by
@@
Second, right-censoring at $M>0$ of this lower-truncated claim leads to a point mass in $M$, resulting in the density $f$ of the lower-truncated and right-censored claim $Y$
@@
assuming that the response variable $Y$ is absolutely continuous on $(0,M)$ with density $f_{\theta}(y)$, having a point mass $1-F_{\theta}(M_{-})=1-\lim_{y\uparrow M}F_{\theta}(y)$ in $M$, and with (unknown) model parameter $\theta$. Fitting such a model with MLE can be difficult because we need an analytically tractable form for both the density $f_{\theta}(\cdot)$ and its distribution function $F_{\theta}(\cdot)$, see (1.3). This is not the case, e.g., in the lower-truncated and right-censored gamma model given in (1.2). Therefore, in such cases, one either needs to rely on numerical integration of the density (which can be computationally demanding, e.g., when performing a regression with fixed covariates) or one uses a version of the Expectation-Maximization (EM) algorithm by interpreting the lower-truncation and right-censoring as a missing information problem; we refer to Verbelen et al. [16], Fung et al. [5] and Sections 6.4.2 and 6.4.3 in Wüthrich–Merz [17]. However, also this EM algorithm approach has its drawbacks as it requires tractability of conditional tail expectations and reasonable dispersion estimates in multi-dimensional parameter settings. These two side constraints lead to further restrictions on the class of solvable models, e.g., these problems can only be solved for a very small number of models within the class of Tweedie’s models [15], namely, for the Tweedie’s models stated in Theorem 3 of Blæsild–Jensen [3]; we also refer to Landsman–Valdez [6]. In a series of papers, Poudyal [9, 10] and Poudyal–Brazauskas [11] consider trimmed and/or winsorized methods of moments estimators for truncated and/or censored data; in statistics, truncation is also called trimming and censoring winsorizing. In these papers, trimming and winsorizing is also shown to be a useful method of robustifying moment estimation under extreme claims.
@@
The goal of this paper is first to study the properties of these MBBEFD densities and to extend it to a bigger class of models that will be called the Bernegger class. Our contribution is to show that the Bernegger class is a rich family of distributions including monotonically decreasing densities, unimodal densities and monotonically increasing densities, and our extension provides new families of lower-truncated and right-censored random variables that allow for skewness in the absolutely continuous part of the distribution. This is of particular interest because unimodal skewed densities are suited for modeling lower-truncated and right-censored claims in general insurance since their empirical density roughly looks like the one given in Figure 1. In particular, the distributions of lower-truncated and right-censored exponential and logistic random variables belong to the Bernegger class.
@@
Surprisingly, the class of MBBEFD densities of Bernegger [2] has only entered the reinsurance literature; see, e.g., Parodi–Watson [8], Abramson [1], Riegel [13], Chapter 21 of Parodi [7], and the R [12] package mbbefd of Dutang et al. [4, 14]. Popular examples in reinsurance pricing are the so-called Swiss Re and Lloyd’s exposure curves that are special cases of MBBEFD exposure curves; see Bernegger [2]. However, in this reinsurance pricing literature, one mainly focuses on exposure curves and not on the resulting densities nor on their properties. We show that the most popular choices from reinsurance lead to monotonically decreasing densities, whereas we are mainly interested into the unimodal case, as this is the common situation in general insurance pricing. By fitting a real dataset consisting of private property insurance claims, we introduce a couple of explicit models belonging to the Bernegger class that allow for unimodal and skewed densities.
Finally, we consider the situation where the insurer is interested in understanding how a change in the deductible or the maximal cover affects the expected claim size. For this, we first emphasize that, typically, the insurer only observes the lower-truncated and right-censored claim $Y$ given in (1.1). That is, for statistical modeling, neither are the claims below the deductible $d$ known, nor are the exact claim sizes above the maximal cover $M$ known. Thus, we can only fit a lower-truncated and right-censored density to observations, e.g., of type (1.2). In general, this does not allow us to extrapolate below the lower-truncation point and above the right-censoring point since there are infinitely many candidates for extrapolation. Under some assumptions on the original density and its support, we can at best smoothly extrapolate, e.g., as the dotted lines in Figure 1 suggest, but the true model could also look completely different, as $Y$ does not reveal any information about the claims being outside of its observed support, except for the proportion of claims exceeding the maximal cover. Therefore, we can only perform the opposite operation of either increasing the deductible or decreasing the maximal cover, and we will show that the Bernegger class is closed under these transformations.
@@
Organization. This manuscript is organized as follows. In the next section, we state the necessary properties that any exposure curve has to fulfill in order to describe distribution functions allowing to model lower-truncated and right-censored claims. In Section 3, we start from the MBBEFD class of distributions of Bernegger [2] by stating some of its properties, and then we extend it to a richer family of distributions, the Bernegger class. In Section 4, we introduce a subclass of the Bernegger class that incorporates the logistic distribution as well as the MBBEFD class of distributions, whereas in Section 5, we treat another subclass of distributions that includes the lower-truncated and right-censored exponential distribution. In Section 6, we use a real dataset of lower-truncated and right-censored claims in order to compare the performance of the gamma and the log-normal model to five examples belonging to the Bernegger class, fitting all these models using maximum likelihood estimation (MLE). Finally, in Section 7, we consider the influence of a change in the deductible or the maximal cover on the observed claims distribution. The last section concludes this work. All mathematical proofs and parameters of the fitted models are provided in the appendix.
@@ 2.1 From Exposure curves to distributions
In reinsurance claims modeling, one often works with exposure curves instead of distribution functions. Assume we have a positively supported response variable $Y\sim F_{Y}$, and assume that the maximal possible loss (MPL) is given by $M>0$, i.e., $0<Y\leq M$, ${\mathbb{P}}$-a.s. We define the normalized loss $Z=Y/M$. Denote the distribution function of the normalized loss $Z$ by $F_{Z}$, being supported in $(0,1]$. The exposure curve of a normalized loss $Z\sim F_{Z}$ is defined by
@@
for $z\in[0,1]$; see Bernegger [2]. This exposure curve $G:[0,1]\to[0,1]$ is non-decreasing, concave, and satisfies the property $G^{\prime}(0)>0$, as well as the normalizations $G(0)=0$ and $G(1)=1$; for examples see Figure 2 (lhs), below. In the last integral, a change of variable $s\in[0,1]\mapsto t=sM\in[0,M]$ gives us
@@
the latter being the exposure curve of $Y$ on $[0,M]$. Thus, we can equally work with the responses $Y$ and $Z$, but the normalized losses $Z$ will have the advantage that they live on the common unit interval $[0,1]$. Now, let us take the opposite view and characterize the distribution function of a random variable $Z\sim F_{Z}$ obtained from a function $G:[0,1]\rightarrow\mathbb{R}$ satisfying the same properties of an exposure curve. Under the assumptions of the next theorem, this distribution $F_{Z}$ leads to an absolutely continuous density on $[0,1)$ and a point mass in 1.
@@
for $z\in[0,1)$, and a point mass in 1 given by
@@
As seen previously, if we start from any such function $G$, we can derive a distribution whose density is absolutely continuous and of closed form on $[0,1)$, with a point mass in 1 and a mean that are of closed form too, i.e., we have a class of models that has fully tractable mean, density and point mass, which is suitable to model right-censored claims. Moreover, if $G^{\prime\prime}(0)<0$, which implies $f_{Z}(0)>0$, it includes lower-truncation in the sense that the density of a lower-truncated random variable $Z$ is positive in zero, see Figure 1. The next result shows that a linear combination of exposure curves allows us to define a mixture of their respective associated distribution functions.
@@
Let $(\alpha_{i})_{i=1}^{n}$ be non-negative weights adding up to 1 and let $(G_{i})_{i=1}^{n}$ be exposure curves leading to distributions functions $(F_{i})_{i=1}^{n}$, densities $(f_{i})_{i=1}^{n}$, and point masses in 1 equal to $(p_{i})_{i=1}^{n}$, as in Theorem 2.1. The convex combination
@@
for $z\in[0,1]$, is again an exposure curve allowing to define the distribution function of a random variable $Z\sim F_{Z}$ given by
@@ 2.2 Flexibility of the point mass in the right-censoring point
The point mass $p$ in 1 is automatically determined by formula (2.3). Often, one may require more modeling flexibility in the choice of this point mass, while still retaining the tractability of the density and the mean as it was shown in Theorem 2.1. A simple way to do so connects to so-called one-inflated distributions; we refer to Dutang et al. [4]. In the case of Theorem 2.1, this can easily be achieved. The next corollary shows, how we can obtain them by looking at the conditional density of the random variable $Z_{0}\stackrel{{\scriptstyle\rm(d)}}{{=}}Z|_{\{Z<1\}}$, which corresponds to a lower- and upper-truncated random variable.
@@
Let $G:[0,1]\rightarrow\mathbb{R}$ be an exposure curve, we receive an absolutely continuous density on $[0,1)$
@@
This density $f_{0}$ integrates to 1 and provides the mean for the random variable $Z_{0}\sim f_{0}$
@@
We can now add a point mass $q\in(0,1)$ in 1 to this density. This adds one more parameter to the model, giving us a mixture distribution between an absolutely continuous part $f_{0}$ on $[0,1)$ and a point mass in 1. We have the following corollary.
@@
Note that the transformation achieved in Corollary 2.5 can also be obtained using Lemma 2.3. Indeed, let us consider the exposure curve
@@
for $z\in[0,1]$, $w\in[0,1]$, and where $G$ is an exposure curve. Using Theorem 2.1, if we denote by $F$ the distribution function obtained from the exposure curve $G$, by $f$ the absolutely continuous density on $[0,1)$ and by $p$ the point mass in 1, we can characterize the distribution function $\tilde{F}$, the absolutely continuous density on $[0,1)$ $\tilde{f}$, and the point mass $\tilde{p}$ obtained from the exposure curve $\tilde{G}$, respectively, using
@@
Remark 2.6.
Note that similarly to Corollary 2.5, one could also add a point mass in 00, thus modeling lower- and right-censored insurance claims with tractable densities, means and point masses.
In what follows, the goal will be to introduce examples of exposure curves that are useful to model lower-truncated and right-censored insurance losses. For this, we start by studying the explicit family of exposure curves introduced by Bernegger [2].
@@ 3.1 The class of MBBEFD exposure curves and densities
The MBBEFD class of Bernegger [2] selects an explicit family of exposure curves. This family is characterized through two parameters $g\geq 1$ and $b\geq 0$, and it is given as follows for $z\in[0,1]$,
@@
${\mathbb{E}}_{b,g}[Z]=\left\{\begin{array}[]{ll}\frac{\log(g)}{g-1}&\text{ for$g>1$ and $b=1$,}\\ \frac{b-1}{\log(b)}&\text{ for $g>1$ and $bg=1$,}\\ \frac{b-1}{\log(b)}\frac{\log(bg)}{bg-1}&\text{ for $g>1$, $b>0$, $b\neq 1$ and $bg\neq 1$.}\end{array}\right.$
@@
For $c=1.5,2,3,4$, one obtains the Swiss Re exposure curves, and for $c=5$, the Lloyd’s exposure curve, these are illustrated in Figure 2 (lhs). The right-hand side of this figure shows the resulting densities $f_{b(c),g(c)}(\cdot)$, and we remark that all of the considered MBBEFD densities are monotonically decreasing on $[0,1]$. $\blacksquare$